# Theory of Unsupervised Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Estimating Background Covariances via Super-Resolution

### v4

@@ 5 Methods
5.1 Fluid simulation
An idealized barotropic ocean jet, proposed by David et al. [51], was simulated in a two-dimensional periodic channel, using the same configuration as in Yasuda and Onishi [35]. The experimental setup is briefly explained here. See the above references for further details.
The governing equations in the periodic channel $(x,y)$ are as follows [51]:
@@ 5.1 Fluid simulation
where $x$ is the channel-wise direction and $y$ is the transverse direction ($x\in[0,2\pi]$ and $y\in[0,\pi]$). Equation (19a) describes the evolution of vorticity $\omega$ over time $t$. The left-hand side represents the advection of $\omega$ and the beta effect [52], while the right-hand side comprises linear drag, hyperviscosity, and forcing due to zonal wind stress $\tau(y)$. Equations (19b) and (19c) define the stream function $\psi$ and the velocity $\bm{u}$, respectively.
@@
where $\tau_{0}=0.3$, $y_{0}=\pi/2$, $\delta=0.4$, and $c$ was adjusted such that the integral of $\tau(y)$ was zero. This setup falls within the parameter regime of mixing barriers with strong eddies [51], where coherent vortices persist in a statistically steady state (Section 6.1).
The initial condition was a zonal jet superimposed with random perturbations. The zonal jet adopts the same shape as $\tau(y)$ and satisfies the Rayleigh-Kuo inflection-point criterion [53], which is a necessary condition for barotropic instability. Perturbations were added to the vorticity field for each wavenumber component, with phases and amplitudes drawn from Gaussian distributions. The data for deep learning were generated by varying these random perturbations.
@@
Table 1: Fluid simulation configurations at two different spatial resolutions: LR and HR. The cutoff wavenumber is the maximum wavenumber magnitude in the CFD model, which was determined by the specific condition to avoid aliasing errors in the pseudo-spectral method (i.e., the 2/3-rule) [54].
@@ 5.2 Synthetic observations
Synthetic observations were generated from the HR simulations. We consider an idealized point-wise observation network that imitates moored buoys in the ocean or balloon-borne radiosondes in the atmosphere. Observation points were subsampled at constant spatial intervals of $8\times 8$ grid points, which means that 1.56% ($=1/8^{2}$) of grid points were observed. To demonstrate that the SRDA can address changes in observation points, the subsampled points were varied in time by random shifts. For the unobserved points, Not a Numbers (NaNs) were assigned and later replaced with missing values in the preprocessing (Section 5.3). To emulate measurement errors, spatially independent Gaussian noise was added to the subsampled vorticity. The noise has a mean of 0 and a standard deviation of 0.1. This standard deviation is approximately 5% of the spatio-temporal average of the absolute values of vorticity.
5.3 Neural networks
@@
We outline the network architectures of an SR model and a CVAE for the sequential SRDA. Hyperparameters are provided in the Appendix, and the full implementation is available at our Zenodo repository (see Data Availability Statement). PyTorch 1.11.0 [55] was used for the implementation.
@@
The SR model architecture is shown in the top panel of Fig. 7. This model resembles the neural network for SR known as VDSR, proposed by Kim et al. [56]. The input is the two-dimensional vorticity field obtained from the LR fluid model, namely the LR forecast state, which is super-resolved to the HR. First, this input is nonlinearly transformed through a stack of convolution and ReLU layers. Upsampling is then executed by two blocks that include pixel shuffle layers [57]. Each block doubles the size of each dimension, resulting in a fourfold increase in resolution. Additionally, a skip connection [58] is employed, whereby the LR input, resized by bicubic interpolation, is added to the final output. This skip connection allows the SR model to capture the discrepancies between the LR input and the HR output.
@@
Figure 7: Network architectures of the SR model and the CVAE. The latter consists of an encoder and a decoder. The label “Res Block” refers to a residual block, and “Res Block x6” refers to the stacking of six Res Blocks. As in the typical implementation of VAEs [7], the output of variance is regarded as $\ln V$ because neural networks can generally output both positive and negative values. Consequently, $\bm{a}_{\text{HR}}+\bm{\varepsilon}^{\text{T}}\exp(1/2\ln V)$ provides a realization of the true state $\bm{\hat{z}}$ in Eq. (14a). All inputs and outputs consist only of vorticity (i.e., 1 channel).
@@
The CVAE consists of an encoder and a decoder (Fig. 7). The encoder performs SRDA and integrates the inputs, an LR forecast and an HR observation, into the HR analysis. The HR observation is first downsampled through convolution to match its resolution with the LR (downsampled by a factor of 4). The LR forecast and the downsampled observation are then concatenated and passed through residual blocks to obtain the LR analysis $\bm{a}_{\text{LR}}$. This $\bm{a}_{\text{LR}}$ is super-resolved by the trained SR model to the HR analysis $\bm{a}_{\text{HR}}$. The variance for $\bm{a}_{\text{HR}}$ [$V$ in Eq. (12)] is also estimated by the encoder. Although $\bm{a}_{\text{HR}}$ and $V$ can be treated independently in the encoder (Section 4.5), we compute $V$ from $\bm{a}_{\text{LR}}$ to reduce the number of parameters, using the feature extractor in the trained SR model (i.e., Feature Extractor-1 in Fig. 7).
@@
The decoder acts as an observation operator and reconstructs the observations from $\bm{\hat{z}}$. This operation comprises convolution and Leaky ReLU [59]. We confirmed that the decoder can be replaced with a subsampling operation (not a neural network), though this replacement slightly decreases the inference accuracy. This result is due to the experimental setup where the synthetic observations were generated by subsampling (Section 5.2).
@@
The clipping function, ${\text{clip}}_{[0,1]}(z)=\min\{1,\max\{0,z\}\}$ ($z\in\mathbb{R}$), restricts the value range to the interval $[0,1]$. The parameters $m_{\omega}$ and $s_{\omega}$ are determined such that 99.9% of the vorticity values fall within the $[0,1]$ range. After applying Eq. (21), zero is regarded as a missing value and is assigned to the HR grid points without observations. This clear interpretation of zero facilitates feature extraction from the input data [60].
5.4 Training and testing methods for the neural networks
5.4.1 Training method
We adopted an offline strategy in which training data were pre-generated (Fig. 8a). The training data consist of pairs of HR observations and LR forecasts, as discussed in Sections 4.4 and 4.5. The initial conditions for the LR fluid model were obtained by applying a low-pass filter to the HR simulation results in the wavenumber domain. This filtering completely removes all high-wavenumber components using discrete Fourier transform. After time integration over $\Delta T$ ($=1$), the resultant LR forecasts were used in training. This $\Delta T$ equals the assimilation interval, which is a reasonable timescale because $\Delta T=1$ is approximately half of the advection time of the zonal jet. The corresponding observations were obtained by subsampling. Figure 8a denotes pairs of the LR forecasts and HR observations as stars. By varying the random initial perturbations (Section 5.1), we conducted 4,500 simulations; 70% were used as training data and the remaining 30% as validation data for hyperparameter tuning.
@@ 5.4.1 Training method
In principle, HR simulations are not required because the training data include LR forecasts rather than HR results. These LR forecasts are obtained by numerical integration in which the analysis states are used as initial conditions. In offline training, however, such an analysis is unavailable because the CVAE is not trained yet. Thus, the filtered HR results were used as the initial conditions. In realistic applications, these initial conditions can be obtained from existing assimilated data [61, 62], or online training would be effective [63, 64], where training and data generation are performed alternately or in parallel.
@@
The SR model was trained, followed by training the CVAE. For both trainings, the Adam optimizer [65] was used with a learning rate of $1\times 10^{-4}$ and a mini-batch size of 128. Each training was terminated using early stopping with a patience parameter of 50 epochs. During the CVAE training, we fixed all parameters in the SR model. The loss functions of the SR model and the CVAEs are $l_{\text{SR}}$ in Eq. (18) and $l_{\text{SRDA}}$ in Eq. (13), respectively. The parameters in $l_{\text{SRDA}}$ were set to $r=1.2\times 10^{-5}$ and $b=1.0\times 10^{-3}$. The value of $r$ was determined by the variance of the measurement errors, while $b$ was determined through grid search to balance the reconstruction error and the KL divergence in the ELBO. Without this balancing, the performance of the CVAE would deteriorate. This phenomenon frequently occurs in VAEs [66, 67], and the loss balancing can be optimized during training [68]. For simplicity, however, this advanced technique was not employed in our study.
5.4.2 Testing method
@@
The sequential SRDA was tested using the method described in Section 4.6. Figure 8b shows a schematic of the test phase. The only difference between Figs. 6 and 8b is that observations are generated by subsampling the HR simulation results. Due to this difference, the HR results are regarded as the ground truth, and the SRDA is evaluated by comparing its inferences to the ground truth. We newly conducted 500 HR fluid simulations. For each case, the LR fluid model was initialized with the low-pass filtered HR vorticity at $t=0$; the SRDA was subsequently conducted until $t=20$ with the assimilation interval $\Delta T=1$. Without SRDA, the vorticity evolution would significantly differ from the ground truth, not only due to differences in the initial condition (i.e., all high-wavenumber components are completely removed) but also due to the difference in the fluid model resolution (i.e., LR versus HR).
@@
Two test metrics were used to assess pixel-wise accuracy and pattern consistency, both of which measure deviations from the ground truth. These two types of metrics are frequently used in SR studies [29] because they complement each other. The pixel-wise accuracy quantifies localized differences but may overestimate discrepancies when spatial patterns are correctly inferred with a slight shift. The pattern consistency mitigates this effect by evaluating structural similarity between inferred and ground-truth patterns, although it may overlook significant pixel-wise deviations.
@@
The mean structural similarity index measure (MSSIM) loss evaluates the consistency of spatial patterns [69]:
@@
where ${\text{C}_{1}}=0.01$, ${\text{C}_{2}}=0.03$, $\mu_{i}$ and $\sigma_{i}^{2}$ are the mean and variance of the ground truth, respectively, $\hat{\mu}_{i}$ and $\hat{\sigma}_{i}^{2}$ are the corresponding quantities for the inference, and $\gamma_{i}$ is the covariance between the ground truth and the inference. These variables in the summation are calculated locally in space using a Gaussian filter. The MSSIM loss takes a value greater than or equal to zero; a smaller value indicates spatial patterns more similar to the ground truth. A detailed discussion of MSSIM can be found in Wang et al. [69].
The MAE ratio and MSSIM loss were calculated at each time step in each test simulation. The values averaged across all test simulations are referred to by the same names. Both MAE ratio and MSSIM loss are collectively called the test errors.
5.5 Baseline model using Ensemble Kalman Filter (EnKF)
As a baseline for comparison with the SRDA, we employed an ensemble Kalman filter (EnKF) [70] for the LR fluid model, using the perturbed observation method [71]. The choice of EnKF was motivated by its ability to estimate uncertainty through ensemble simulations. The uncertainties estimated by the EnKF and SRDA are compared in Section 6.4.
Assimilation was conducted in the HR space after the LR forecast was super-resolved to the HR using bicubic interpolation, with the assimilation interval $\Delta T$ ($=1$) being the same as in the SRDA. Barthélémy et al. [34] argued that performing the EnKF in the HR space can yield more accurate inferences than those in the LR space when observations are defined at HR grid points. Furthermore, this approach clarifies the comparison with the SRDA because both EnKF and SRDA employ the LR fluid model and perform DA in the HR space.
The background error covariance matrices were spatially localized using a function proposed by Gaspari and Cohn [72]. These error covariances were inflated by adding Gaussian noise to the analysis before the forecast process [73]. This Gaussian noise incorporates spatial correlation by estimating the covariance from the training data at each time step. The initial perturbations for creating ensemble members were also drawn from this Gaussian distribution.
Hyperparameters for the EnKF were adjusted using the training data to minimize the MAE. Specifically, the following parameters were tuned: the number of ensemble members (set to 300 after tuning), the amplitude of the initial perturbation for generating ensemble members, the amplitude of the additive inflation, the localization radius, and the amplitude for perturbing observations. The full implementation of the EnKF is also available at our Zenodo repository (see Data Availability Statement).
6 Results and discussion
6.1 Time evolution of vorticity in the jet system

### v5

@@ V Methods
V.1 Fluid simulation
An idealized barotropic ocean jet, proposed by David et al.,David, Marshall, and Zanna (2017) was simulated in a two-dimensional periodic channel, using the same configuration as in Yasuda and Onishi.Yasuda and Onishi (2023) The experimental setup is briefly explained here. See the above references for further details.
The governing equations in the periodic channel $(x,y)$ are as follows:
@@ V.1 Fluid simulation
where $x$ is the channel-wise direction and $y$ is the transverse direction ($x\in[0,2\pi]$ and $y\in[0,\pi]$). Equation (19a) describes the evolution of vorticity $\omega$ over time $t$. The left-hand side represents the advection of $\omega$ and the beta effect,Vallis (2017a) while the right-hand side comprises linear drag, hyperviscosity, and forcing due to zonal wind stress $\tau(y)$. Equations (19b) and (19c) define the stream function $\psi$ and the velocity $\bm{u}$, respectively. All variables and parameters are non-dimensionalized following David et al.David, Marshall, and Zanna (2017)
@@
where $\tau_{0}=0.3$, $y_{0}=\pi/2$, $\delta=0.4$, and $c$ was adjusted such that the integral of $\tau(y)$ was zero. This setup falls within the parameter regime of mixing barriers with strong eddies,David, Marshall, and Zanna (2017) where coherent vortices persist in a statistically steady state (Section VI.1).
The initial condition was a zonal jet superimposed with random perturbations. The zonal jet adopts the same shape as $\tau(y)$ and satisfies the Rayleigh-Kuo inflection-point criterion,Vallis (2017b) which is a necessary condition for barotropic instability. Perturbations were added to the vorticity field for each wavenumber component, with phases and amplitudes drawn from Gaussian distributions. The data for deep learning were generated by varying these random perturbations.
@@
Table 1: Fluid simulation configurations at two different spatial resolutions: LR and HR. The cutoff wavenumber is the maximum wavenumber magnitude in the CFD model, which was determined by the specific condition to avoid aliasing errors in the pseudo-spectral method (i.e., the 2/3-rule).Canuto et al. (1988)
@@ V.2 Synthetic observations
Synthetic observations were generated from the HR simulations. We consider an idealized point-wise observation network that imitates moored buoys in the ocean or balloon-borne radiosondes in the atmosphere. Observation points were subsampled at constant spatial intervals of $8\times 8$ grid points, which means that 1.56% ($=1/8^{2}$) of grid points were observed. To demonstrate that the SRDA can address changes in observation points, the subsampled points were varied in time by random shifts. For the unobserved points, Not a Numbers (NaNs) were assigned and later replaced with missing values in the preprocessing (Section V.3). To emulate measurement errors, spatially independent Gaussian noise was added to the subsampled vorticity. This noise has a mean of 0 and a standard deviation of 0.1. This standard deviation is approximately 5% of the spatio-temporal average of the absolute values of vorticity.
V.3 Neural networks
@@
We outline the network architectures of an SR model and a CVAE for the sequential SRDA. Hyperparameters are provided in the Appendix, and the full implementation is available at our Zenodo repositoryYasuda and Onishi (2024) (see Data Availability Statement). PyTorch 1.11.0Paszke et al. (2019) was used for the implementation.
@@
The SR model architecture is shown in the top panel of Fig. 7. This model resembles the neural network for SR known as VDSR, proposed by Kim et al.Kim, Lee, and Lee (2016) The input is the two-dimensional vorticity field obtained from the LR fluid model, namely the LR forecast state, which is super-resolved to the HR. First, this input is nonlinearly transformed through a stack of convolution and ReLU layers. Upsampling is then executed by two blocks that include pixel shuffle layers.Shi et al. (2016) Each block doubles the size of each dimension, resulting in a fourfold increase in resolution. Additionally, a skip connectionHe et al. (2016) is employed, whereby the LR input, resized by bicubic interpolation, is added to the final output. This skip connection allows the SR model to capture the discrepancies between the LR input and the HR output.
@@
Figure 7: Network architectures of the SR model and the CVAE. The latter consists of an encoder and a decoder. The label “Res Block” refers to a residual block, and “Res Block x6” refers to the stacking of six Res Blocks. As in the typical implementation of VAEs,Kingma and Welling (2019) the output of variance is regarded as $\ln V$ because neural networks can generally output both positive and negative values. Consequently, $\bm{a}_{\text{HR}}+\bm{\varepsilon}^{\text{T}}\exp(1/2\log v)$ provides a realization of the true state $\bm{\hat{z}}$ in Eq. (14a). All inputs and outputs consist only of vorticity (i.e., 1 channel).
@@
The CVAE consists of an encoder and a decoder (Fig. 7). The encoder performs SRDA and integrates the two inputs, an LR forecast and an HR observation, into the HR analysis. The HR observation is first downsampled through convolution to match its resolution with the LR (downsampled by a factor of 4). The effectiveness of separating forecast and observation inputs is shown in our previous study.Yasuda and Onishi (2023) The LR forecast and the downsampled observation are then concatenated and passed through residual blocks to obtain the LR analysis $\bm{a}_{\text{LR}}$. This $\bm{a}_{\text{LR}}$ is super-resolved by the trained SR model to the HR analysis $\bm{a}_{\text{HR}}$. The variance for $\bm{a}_{\text{HR}}$ [$V$ in Eq. (12)] is also estimated by the encoder. Although $\bm{a}_{\text{HR}}$ and $V$ can be treated independently in the encoder (Section IV.5), we compute $V$ from $\bm{a}_{\text{LR}}$ to reduce the number of parameters, using the feature extractor in the trained SR model (i.e., Feature Extractor-1 in Fig. 7).
@@
The decoder acts as an observation operator and reconstructs the observations from $\bm{\hat{z}}$. This operation comprises convolution and Leaky ReLU.Maas et al. (2013) We confirmed that the decoder can be replaced with a subsampling operation (not a neural network), though this replacement slightly decreases the inference accuracy. This result is due to the experimental setup where the synthetic observations were generated by subsampling (Section V.2).
@@
The clipping function, ${\text{clip}}_{[0,1]}(z)=\min\{1,\max\{0,z\}\}$ ($z\in\mathbb{R}$), restricts the value range to the interval $[0,1]$. The parameters $m_{\omega}$ and $s_{\omega}$ are determined such that 99.9% of the vorticity values fall within the $[0,1]$ range. After applying Eq. (21), zero is regarded as a missing value and is assigned to the HR grid points without observations. This clear interpretation of zero facilitates feature extraction from the input data.Yasuda, Onishi, and Matsuda (2023)
V.4 Training and testing methods for the neural networks
V.4.1 Training method
We adopted an offline strategy in which training data were pre-generated [Fig. 8(a)]. The training data consist of pairs of HR observations and LR forecasts, as discussed in Sections IV.4 and IV.5. The initial conditions for the LR fluid model were obtained by applying a low-pass filter to the HR simulation results in the wavenumber domain. This filtering completely removes all high-wavenumber components using discrete Fourier transform. After time integration over $\Delta T$ ($=1$), the resultant LR forecasts were used in training. This $\Delta T$ equals the assimilation interval, which is a reasonable timescale because $\Delta T=1$ is approximately half of the advection time of the zonal jet. The corresponding observations were obtained by subsampling. Figure 8(a) denotes pairs of the LR forecasts and HR observations as stars. By varying the random initial perturbations (Section V.1), we conducted 4,500 simulations; 70% were used as training data and the remaining 30% as validation data for hyperparameter tuning.
@@ V.4.1 Training method
In principle, HR simulations are not required because the training data include LR forecasts rather than HR results. These LR forecasts are obtained by numerical integration in which the analysis states are used as initial conditions. In offline training, however, such an analysis is unavailable because the CVAE is not trained yet. Thus, the filtered HR results were used as the initial conditions. In realistic applications, these initial conditions can be obtained from existing assimilated data,Hersbach et al. (2020); KOSAKA et al. (2024) or online training would be effective,Sahoo et al. (2018); Rasp (2020) where training and data generation are performed alternately or in parallel.
@@
The SR model was trained, followed by training the CVAE. For both trainings, the Adam optimizerKingma and Ba (2015) was used with a learning rate of $1\times 10^{-4}$ and a mini-batch size of 128. Each training was terminated using early stopping with a patience parameter of 50 epochs. During the CVAE training, we fixed all parameters in the SR model. Note that VAEs typically show stable training and inference characteristicsKingma and Welling (2014) and we confirmed this behavior in our experiments (not shown). The loss functions of the SR model and the CVAEs are $l_{\text{SR}}$ in Eq. (18) and $l_{\text{SRDA}}$ in Eq. (13), respectively. The parameters in $l_{\text{SRDA}}$ were set to $r=1.2\times 10^{-5}$ and $b=1.0\times 10^{-3}$. The value of $r$ was determined by the variance of the measurement errors, while $b$ was determined through grid search to balance the reconstruction error and the KL divergence in the ELBO. Without this balancing, the performance of the CVAE would deteriorate. This phenomenon frequently occurs in VAEs,Higgins et al. (2017); Alemi et al. (2018) and the loss balancing can be optimized during training.Asperti and Trentin (2020) For simplicity, however, this advanced technique was not employed in our study.
V.4.2 Testing method
@@
The sequential SRDA was tested using the method described in Section IV.6. Figure 8(b) shows a schematic of the test phase. The only difference between Figs. 6 and 8(b) is that observations are generated by subsampling the HR simulation results. Due to this difference, the HR results are regarded as the ground truth, and the SRDA is evaluated by comparing its inferences to the ground truth. We newly conducted 500 HR fluid simulations. For each case, the LR fluid model was initialized with the low-pass filtered HR vorticity at $t=0$; the SRDA was subsequently conducted until $t=20$ with the assimilation interval $\Delta T=1$. Without SRDA, the vorticity evolution would significantly differ from the ground truth, not only due to differences in the initial condition (i.e., all high-wavenumber components are completely removed) but also due to the difference in the fluid model resolution (i.e., LR versus HR).
@@
Two test metrics were used to assess pixel-wise accuracy and pattern consistency, both of which measure deviations from the ground truth. These two types of metrics are frequently used in SR studiesLepcha et al. (2023) because they complement each other. The pixel-wise accuracy quantifies localized differences but may overestimate discrepancies when spatial patterns are correctly inferred with a slight shift. The pattern consistency mitigates this effect by evaluating structural similarity between inferred and ground-truth patterns, although it may overlook significant pixel-wise deviations.
@@
The mean structural similarity index measure (MSSIM) loss evaluates the consistency of spatial patterns:Wang et al. (2004)
@@
where ${\text{C}_{1}}=0.01$, ${\text{C}_{2}}=0.03$, $\mu_{i}$ and $\sigma_{i}^{2}$ are the mean and variance of the ground truth, respectively, $\hat{\mu}_{i}$ and $\hat{\sigma}_{i}^{2}$ are the corresponding quantities for the inference, and $\gamma_{i}$ is the covariance between the ground truth and the inference. These variables in the summation are calculated locally in space using a Gaussian filter. The MSSIM loss takes a value greater than or equal to zero; a smaller value indicates spatial patterns more similar to the ground truth. A detailed discussion of MSSIM can be found in Wang et al.Wang et al. (2004)
The MAE ratio and MSSIM loss were calculated at each time step in each test simulation. The values averaged across all test simulations are referred to by the same names. Both MAE ratio and MSSIM loss are collectively called the test errors. We discuss the results from the test data using these metrics in Section VI.
V.5 Baseline model using Ensemble Kalman Filter (EnKF)
As a baseline for comparison with the SRDA, we employed an ensemble Kalman filter (EnKF)Evensen (1994) for the LR fluid model, using the perturbed observation method.Burgers, van Leeuwen, and Evensen (1998) The choice of EnKF was motivated by its ability to estimate uncertainty through ensemble simulations. The uncertainties estimated by the EnKF and SRDA are compared in Section VI.4.
Assimilation was conducted in the HR space after the LR forecast was super-resolved to the HR using bicubic interpolation, with the assimilation interval $\Delta T$ ($=1$) being the same as in the SRDA. Barthélémy et al.Barthélémy et al. (2022) argued that performing the EnKF in the HR space can yield more accurate inferences than those in the LR space when observations are defined at HR grid points. Furthermore, this approach clarifies the comparison with the SRDA because both EnKF and SRDA employ the LR fluid model and perform DA in the HR space.
The background error covariance matrices were spatially localized using a function proposed by Gaspari and Cohn.Gaspari and Cohn (1999) These error covariances were inflated by adding Gaussian noise to the analysis before the forecast process.Whitaker et al. (2008) This Gaussian noise incorporates spatial correlation by estimating the covariance from the training data at each time step. The initial perturbations for creating ensemble members were also drawn from this Gaussian distribution.
Hyperparameters for the EnKF were adjusted using the training data to minimize the MAE. Specifically, the following parameters were tuned: the number of ensemble members (set to 300 after tuning), the amplitude of the initial perturbation for generating ensemble members, the amplitude of the additive inflation, the localization radius, and the amplitude for perturbing observations. The full implementation of the EnKF is also available at our Zenodo repositoryYasuda and Onishi (2024) (see Data Availability Statement).
VI Results and discussion
VI.1 Time evolution of vorticity in the jet system